Two identical pith balls, each carrying a charge q, are suspended from a common point by two strings of equal length l, Find the mass of each ball if the angle between the strings is 2θ in equilibrium.

Grade:11

2 Answers

Aditi Chauhan
9 years ago
Sol. Electric force = kq^2/(ℓ sin Q + ℓ sin Q)^2 = kq^2/4 ℓ^2 sin^2 So, T Cos θ = ms (For equilibrium) T sin θ = Ef Ef tan θ = Ef/mf ⇒ mg = Ef cot θ = kq^2/4ℓ^2 sin^2θ cot θ = q^2 cot θ/ℓ^2 sin^2θ 16πE base 0 Or m = q^2 cotθ/16πE base 0ℓ^2 Sin^2θg
Abhishek bains
6 years ago
F=mg tanxqq/16πEl^2 sin^2x tanx=m.......?........................... ............................................................. ............